# Chengxuan Qian1,  Huaxiu Yao4, Zhengzhong Tu [^1] footnotemark: 1

### v1

@@ Abstract
The emergence of large Vision Language Models (VLMs) has broadened the scope and capabilities of single-modal Large Language Models (LLMs) by integrating visual modalities, thereby unlocking transformative cross-modal applications in a variety of real-world scenarios. Despite their impressive performance, VLMs are prone to significant hallucinations, particularly in the form of cross-modal inconsistencies. Building on the success of Reinforcement Learning from Human Feedback (RLHF) in aligning LLMs, recent advancements have focused on applying direct preference optimization (DPO) on carefully curated datasets to mitigate these issues. Yet, such approaches typically introduce preference signals in a brute-force manner, neglecting the crucial role of visual information in the alignment process. In this paper, we introduce Re-Align, a novel alignment framework that leverages image retrieval to construct a dual-preference dataset, effectively incorporating both textual and visual preference signals. We further introduce rDPO, an extension of the standard direct preference optimization that incorporates an additional visual preference objective during fine-tuning. Our experimental results demonstrate that Re-Align not only mitigates hallucinations more effectively than previous methods but also yields significant performance gains in general visual question-answering (VQA) tasks. Moreover, we show that Re-Align maintains robustness and scalability across a wide range of VLM sizes and architectures. This work represents a significant step forward in aligning multimodal LLMs, paving the way for more reliable and effective cross-modal applications. We release all the code in https://github.com/taco-group/Re-Align.
@@
Shuo Xing1thanks:  Email: {shuoxing,tzz}@tamu.edu, Yuping Wang2, Peiran Li1, Ruizheng Bai1, Yueqi Wang3, Chengxuan Qian1,  Huaxiu Yao4, Zhengzhong Tu [^1] footnotemark: 1thanks:   Corresponding author. 1Texas A&M University  2University of Michigan  3UIUC  4UNC Chapel Hill
@@ 1 Introduction
The recent emergence of powerful Vision Language Models (VLMs) (Li et al., 2022, 2023a; Liu et al., 2024a; Li et al., 2024b; Meta, 2024; Bai et al., 2023; Wang et al., 2024b; Lu et al., 2024; Wu et al., 2024) has significantly extended the capabilities of Large Language Models (LLMs) (Devlin et al., 2018; Radford et al., 2019; Brown et al., 2020; Team et al., 2023; Roziere et al., 2023; Touvron et al., 2023a, b; Raffel et al., 2020; Yang et al., 2024; Team, 2024) into the visual domain, paving the way for innovative real-world applications that integrate multimodal information (Moor et al., 2023; Li et al., 2024a; Shao et al., 2024; Xing et al., 2024b; Rana et al., 2023; Kim et al., 2024). Despite their promising performance, VLMs remain susceptible to hallucinations—instances where the model produces outputs containing inaccurate or fabricated details about objects, attributes, and the logical relationships inherent in the input image (Rohrbach et al., 2018; Bai et al., 2024). Several factors contribute to this cross-modal inconsistencies, including the separate low-quality or biased training data, imbalanced model architectures, and the disjoint pretraining of the vision encoder and LLM-backbone (Cui et al., 2023; Bai et al., 2024; Zhou et al., 2024a).
To mitigate the hallucinations in VLMs, the Directed Preference Optimization (DPO) techniques have been widely adopted (Deng et al., 2024; Zhou et al., 2024a; Fang et al., 2024; Zhou et al., 2024b; Guo et al., 2024; Chen et al., 2024; Wang et al., 2024c; Yu et al., 2024b; Li et al., 2023b; Wang et al., 2024a). This involves constructing datasets enriched with human preference signals specifically targeting hallucinations, and then finetuning the models using algorithms like Direct Preference Optimization (DPO) (Rafailov et al., 2024). Existing methods generating the preference data by perturbing the ground truth responses (Zhou et al., 2024a) and corrupting the visual inputs/embeddings (Deng et al., 2024; Amirloo et al., 2024) to generate rejected responses or correcting/refining responses to produce chosen responses (Chen et al., 2024; Yu et al., 2023a). While methods based on response refinement yield the most reliable preference signals, they face scalability challenges due to the significant costs of manual correction processes. Conversely, directly corrupting input visual information or ground truth responses is overly simplistic, as this brute-force approach fails to generate plausible and natural hallucinations in a controlled manner. Moreover, during fine-tuning, directly applying DPO may cause the model to overly prioritize language-specific preferences, which potentially leads to suboptimal performance and an increased propensity for hallucinations (Wang et al., 2024a).
@@ 3 Methods
As shown in Figure 2, the process begins with an advanced VLM generating chosen responses from input images from the training set. A selective masking process is then applied, strategically omitting segments associated with objects, attributes, or logical relationships identified in the image. Next, leveraging the retrieved image from the same training dataset and the masked responses, the hallucination-prone VLM is prompted to complete the masked elements, obtaining rejected responses. The generated preference pairs (chosen vs. rejected) are then used to fine-tune the VLM with $\mathcal{L}_{\text{rDPO}}$ (eq. 1), a preference objective that integrates both visual and textual information to penalize hallucinations and reinforce grounded reasoning. Algorithm 1 provides an overview of Re-Align, while the detailed process is explained in the following subsections.

### v2

# Chan-wei Hu1, Chengxuan Qian1, Huaxiu Yao4, Zhengzhong Tu [^1] footnotemark: 1

@@ Abstract
The emergence of large Vision Language Models (VLMs) has broadened the scope and capabilities of single-modal Large Language Models (LLMs) by integrating visual modalities, thereby unlocking transformative cross-modal applications in a variety of real-world scenarios. Despite their impressive performance, VLMs are prone to significant hallucinations, particularly in the form of cross-modal inconsistencies. Building on the success of Reinforcement Learning from Human Feedback (RLHF) in aligning LLMs, recent advancements have focused on applying direct preference optimization (DPO) on carefully curated datasets to mitigate these issues. Yet, such approaches typically introduce preference signals in a brute-force manner, neglecting the crucial role of visual information in the alignment process. In this paper, we introduce Re-Align, a novel alignment framework that leverages image retrieval to construct a dual-preference dataset, effectively incorporating both textual and visual preference signals. We further introduce rDPO, an extension of the standard direct preference optimization that incorporates an additional visual preference objective during fine-tuning. Our experimental results demonstrate that Re-Align not only mitigates hallucinations more effectively than previous methods but also yields significant performance gains in general visual question-answering (VQA) tasks. Moreover, we show that Re-Align maintains robustness and scalability across a wide range of VLM sizes and architectures. This work represents a significant step forward in aligning multimodal LLMs, paving the way for more reliable and effective cross-modal applications.
@@
Shuo Xing1thanks:  Email: {shuoxing,tzz}@tamu.edu, Yuping Wang2, Peiran Li1, Ruizheng Bai1, Yueqi Wang3, Chan-wei Hu1, Chengxuan Qian1, Huaxiu Yao4, Zhengzhong Tu [^1] footnotemark: 1thanks:   Corresponding author. 1Texas A&M University  2University of Michigan  3UIUC  4UNC Chapel Hill
@@ 1 Introduction
The recent emergence of powerful Vision Language Models (VLMs) (Li et al., 2022, 2023a; Liu et al., 2024a; Li et al., 2024b; Meta, 2024; Bai et al., 2023; Wang et al., 2024b; Lu et al., 2024; Wu et al., 2024) has significantly extended the capabilities of Large Language Models (LLMs) (Devlin et al., 2018; Radford et al., 2019; Brown et al., 2020; Team et al., 2023; Roziere et al., 2023; Touvron et al., 2023a, b; Raffel et al., 2020; Yang et al., 2024; Team, 2024) into the visual domain, paving the way for innovative real-world applications that integrate multimodal information (Moor et al., 2023; Li et al., 2024a; Shao et al., 2024; Xing et al., 2024b; Rana et al., 2023; Kim et al., 2024). Despite their promising performance, VLMs remain susceptible to hallucinations—instances where the model produces outputs containing inaccurate or fabricated details about objects, attributes, and the logical relationships inherent in the input image (Rohrbach et al., 2018; Bai et al., 2024). Several factors contribute to this cross-modal inconsistency, including the separate low-quality or biased training data, imbalanced model architectures, and the disjoint pretraining of the vision encoder and LLM-backbone (Cui et al., 2023; Bai et al., 2024; Zhou et al., 2024a).
To mitigate the hallucinations in VLMs, the Directed Preference Optimization (DPO) techniques have been widely adopted (Deng et al., 2024; Zhou et al., 2024a; Fang et al., 2024; Zhou et al., 2024b; Guo et al., 2024; Chen et al., 2024; Wang et al., 2024c; Yu et al., 2024b; Li et al., 2023b; Wang et al., 2024a; Xiao et al., 2025; Xie et al., 2024; Fu et al., 2024). This involves constructing datasets enriched with human preference signals specifically targeting hallucinations, and then finetuning the models using algorithms like Direct Preference Optimization (DPO) (Rafailov et al., 2024). Existing methods generate the preference data by perturbing the ground truth responses (Zhou et al., 2024a) and corrupting the visual inputs/embeddings (Deng et al., 2024; Amirloo et al., 2024) to generate rejected responses or correcting/refining responses to produce chosen responses (Chen et al., 2024; Yu et al., 2023a). While methods based on response refinement yield the most reliable preference signals, they face scalability challenges due to the significant costs of manual correction processes. Conversely, directly corrupting input visual information or ground truth responses is overly simplistic, as this brute-force approach fails to generate plausible and natural hallucinations in a controlled manner. Moreover, during fine-tuning, directly applying DPO may cause the model to overly prioritize language-specific preferences, which potentially leads to suboptimal performance and an increased propensity for hallucinations (Wang et al., 2024a).
@@ 3 Methods
As shown in Figure 2, the process begins with an advanced VLM generating chosen responses from input images from the training set. A selective masking process is then applied, strategically omitting segments associated with objects, attributes, or logical relationships identified in the image. Next, leveraging the retrieved image from the same training dataset and the masked responses, the hallucination-prone VLM is prompted to complete the masked elements, obtaining rejected responses. The generated preference pairs (chosen vs. rejected) are then used to fine-tune the VLM with $\mathcal{L}_{\text{rDPO}}$ (eq. 1), a preference objective that integrates both visual and textual information to penalize hallucinations and reinforce grounded reasoning. Algorithm 1 in Appendix A provides an overview of Re-Align, while the detailed process is explained in the following subsections.
3.1 Preference Generation
Generating high-quality preference data, which includes both accurate ground-truth responses and controlled hallucinated examples, is crucial for effective preference optimization in pre-trained VLMs. Existing methods construct preference data by perturbing ground-truth responses (Zhou et al., 2024a), corrupting visual inputs/embeddings (Deng et al., 2024; Amirloo et al., 2024) to create rejected responses, or refining responses to obtain chosen responses (Chen et al., 2024; Yu et al., 2023a). Refinement produces high-quality preference data but comes at a high cost, whereas direct corruption is more scalable yet tends to generate unrealistic hallucinations and fails to produce plausible, natural ones in a controlled manner. To address these limitations, we introduce a novel image retrieval-based pipeline for preference data construction as shown in Figure 3, which consists of three key stages:
3.2 Preference Optimization
w. LLaVA-RLHF 84.77 84.60 83.40 10.2859 18.7861 38.2418 40.6744 44.6528
w. POVID 88.21 87.16 85.06 10.5495 18.2081 41.5385 40.9302 46.6785
w. SIMA 88.10 87.10 85.03 10.9890 17.6301 43.0549 40.2326 45.2728
w. mDPO 88.17 87.13 85.03 9.8901 18.4971 41.978 40.000 46.1470
w. STIC 89.03 88.20 86.56 12.9670 17.3410 47.2527 34.1860 46.3242
LLaVA-v1.5-7B 66.02 58.18 31.6 50.03 64.1 1510.28 357.85 64.60 3.875
w. LLaVA-RLHF 63.11 56.89 31.8 49.57 60.2 1378.90 282.85 64.39 6
w. POVID 65.98 58.18 31.8 49.80 67.3 1495.91 356.07 64.34 4.375
w. CSR (3Iter) 65.46 57.86 31.6 47.02 68.3 1525.44 365.35 64.08 4.5
w. SIMA 65.83 58.48 32.0 50.04 66.9 1510.33 371.78 64.60 2.75
w. mDPO 67.53 57.90 31.3 50.04 59.0 1510.74 335.71 64.60 4.25
w. Re-Align 68.10 58.55 32.1 50.06 67.7 1511.79 367.50 64.69 1.375
w. STIC 76.42 63.50 47.3 54.21 81.0 1504.91 308.21 69.16 2.625
w. Re-Align 76.47 64.08 48.3 57.27 81.8 1512.09 318.93 69.42 1.25
4 Experiments
We conduct three categories of experiments to empirically validate the effectiveness of our proposed method. First, we evaluate the ability of Re-Align to mitigate hallucinations and improve generalizability across diverse VQA tasks, demonstrating its consistent superiority over baseline approaches and achieving state-of-the-art performance. Next, we examine Re-Align’s effectiveness in aligning VLMs across various model sizes and architectures, including both text-to-image and unified models, where it delivers substantial performance over vanilla models and existing baselines. Finally, we assess the impact of our proposed rDPO objective in preference optimization, showing that it consistently surpasses standard DPO in aligning VLMs and achieving superior results in both halluciation mitigation and general tasks.
4.1 Re-Align for VLMs Alignment
Datasets
Beslines
We compare our method with several widely adopted alignment frameworks for VLMs, including LLaVA-RLHF (Sun et al., 2023), POVID Zhou et al. (2024a), CSR (Zhou et al., 2024b), SIMA (Wang et al., 2024c), STIC (Deng et al., 2024). For more details on these baselines, please refer to the Appendix.
Experimental Setup
We sample 11k images from the LLaVA-Instruct-150K dataset (Liu et al., 2024a) to construct preference data, as illustrated in Figure 3. These images are initially used to generate QA pairs based on image captions and simple VQA tasks using GPT-4o mini (OpenAI, 2024). Furthermore, the images are encoded using clip-vit-large-patch14 (Radford et al., 2021a) to construct the knowledge base for image retrieval. For rejected responses, we use GPT-4o mini to mask the chosen response and all-mpnet-base-v2 (Reimers and Gurevych, 2019) to compute the similarity between the completed masked response and the original chosen response. We use LLaVA-v1.5-7B (Liu et al., 2024a) and LLaVA-v1.6-Mistral-7B (Li et al., 2024b) as our backbone models and perform Re-Align finetuning for $1$ epoch. All evaluations are conducted with a temperature setting of 0, and baseline results are reproduced using open-sourced model weights.
Results
Table 1 shows the performance of Re-Align compared to baseline methods on hallucination benchmarks. Notably, Re-Align achieves the best among the evaluated methods on both POPE and HallusionBench for LLaVA-v1.5-7B (Liu et al., 2024a) and LLaVA-v1.6-Mistral-7B (Li et al., 2024b), highlighting the effectiveness of our approach in mitigating hallucinations of VLMs. As shown in Table 2, Re-Align can provide generally on-par or better performance than the vanilla models and baseline alignment methods on each evaluated general VQA task, ultimately achieving the best overall results. This finding indicates that Re-Align can enhance hallucination mitigation without compromising general performance.
4.2 Scalability and Generalizability
Experimental Setup
  • Unified Models: encoder-decoder architecture that decouples visual encoding for multimodal understanding and generation. We evaluate Re-Align on Janus-Pro-1B (Chen et al., 2025) and Janus-Pro-7B (Chen et al., 2025).
Methods POPEr POPEp POPEa
Janus-Pro-1B 85.46 85.03 84.13
w. Re-Align 87.53↑2.07 87.33↑2.30 85.86↑1.73
Janus-Pro-7B 88.41 87.30 85.70
w. Re-Align 89.73↑1.32 88.37↑1.07 86.27↑0.57
LLaVA-v1.5-7B 88.14 87.23 85.10
w. LLaVA-RLHF 84.77↓3.37 84.60↓2.63 83.40↓0.50
w. POVID 88.21↑0.07 87.16↓0.07 85.06↓0.04
w. CSR (3Iter) 87.83↓0.31 87.00↓0.23 85.00↓0.10
w. SIMA 88.10↓0.04 87.10↓0.13 85.03↓0.07
w. mDPO 88.17↑0.03 87.13↓0.10 85.03↓0.07
w. Re-Align 88.65↑0.51 87.43↑0.20 85.16↑0.06
LLaVA-v1.5-13B 88.07 87.53 85.60
w. CSR (3Iter) 88.38↑0.31 87.90↑0.37 85.46↓0.14
w. SIMA 88.04↓0.03 87.40↓0.13 85.40↓0.20
w. HSA-DPO 85.01↓3.06 85.00↓2.53 83.86↓1.74
w. Re-Align 90.03↑1.96 89.20↑1.30 86.20↑0.74
LLaVA-v1.6-     Vicuna-7B 88.52 87.63 86.36
w. Re-Align 88.94↑0.42 88.03↑0.40 86.63↑0.27
LLaVA-v1.6-     Vicuna-13B 88.24 87.70 86.43
w. Re-Align 88.79↑0.55 88.10↑0.40 86.60↑0.17
Table 3: Impact of Re-Align across various model scales on POPE.
Results
4.3 Ablation Study
In this section, we conduct a comprehensive ablation study to explore how the data curation framework and design of the objective function affect the Re-Align’ performance. The experimental setup follows the same setting as VLMs alignment experiments.
Dataset
Due to budget constraints and the need for reproducibility, we have excluded benchmarks that require evaluation by GPT-4 (Achiam et al., 2023). Instead, we focus on the following tasks: ScienceQA (Lu et al., 2022), TextVQA (Singh et al., 2019), and POPE (Li et al., 2023c).
$\tau$ SQA TextVQA POPEr POPEp POPEa
0.85 67.04 57.31 88.96 87.83 85.06
0.90 67.75 57.68 88.83 87.66 84.93
0.95 68.10 58.55 88.65 87.43 85.16
Table 4: Impact of similarity threshold $\tau$ for generating the rejected responses in Re-Align across general and hallucination benchmarks for VLMs, and comparisons with baselines.
Similarity Threshold $\tau$
In Re-Align, we set the similarity threshold $\tau$ to 0.95, which acts as an upper bound on the cosine similarity between the chosen response and the generated rejected response. As illustrated in Table 4, decreasing the threshold $\tau$ results in a stronger preference signal, leading to improved performance in mitigating hallucinations. However, this comes at the cost of reduced performance in general VQA. Among the evaluated configurations, setting $\tau=0.95$ offers the best trade-off, effectively reducing hallucinations while maintaining strong performance across VQA benchmarks.
Masking Strategy
In data curation, we generate preference data by inducing hallucinations at the segment level. To further investigate the impact of finer-grained perturbations, we conduct experiments using sentence-level masking. As shown in Table 5, using a sentence-level masking strategy, Re-Align still demonstrates significant improvement in reducing hallucination in VLMs. However, this approach leads to a slight drop in performance on general VQA tasks. More discussions on the masking strategy can be found in Appendix F.
Masking Strategy SQA TextVQA POPEr POPEp POPEa
sentence-level 67.58 57.77 88.56 87.60 84.90
segment-level 68.10 58.55 88.65 87.43 85.16
Table 5: Impact of masking strategy across general and hallucination benchmarks for VLMs, and comparisons with baselines.
Design of Loss Function
In Re-Align, we assign equal weights to the DPO and vDPO objectives in the combined loss function, i.e., $\mathcal{L}_{\text{rDPO}}=\mathcal{L}_{\text{DPO}}+\mathcal{L}_{\text{vDPO}}$. To better understand the impact of this design of loss function, we generalize the loss function to $\mathcal{L}_{\text{DPO}}+w_{v}\mathcal{L}_{\text{vDPO}}$, where $w_{v}$ controls the contribution of the visual component, and conduct experiments with different values of $w_{v}$ to analyze the trade-offs and identify the optimal balance between textual and visual preference signals. As shown in Table 6, incorporating the $\mathcal{L}_{\text{vDPO}}$ objective significantly enhances VLM performance on hallucination benchmarks. In general, when combined with the standard $\mathcal{L}_{\text{DPO}}$ objective, increasing the weight of $\mathcal{L}_{\text{vDPO}}$ tends to yield better overall performance. Notably, the equally-combined objective $\mathcal{L}_{\text{rDPO}}$ achieves the best balance between reducing hallucinations and maintaining strong performance on general VQA benchmarks, highlighting its effectiveness as a robust training strategy.
$w_{v}$ SQA TextVQA POPEr POPEp POPEa
0.0 (DPO) 66.26 58.24 88.18 87.30 85.23
0.25 67.08 58.34 88.93 87.53 84.90
0.50 67.08 58.34 88.93 87.53 84.90
0.75 67.08 58.34 88.93 87.53 84.90
1.0 (rDPO) 68.10 58.55 88.65 87.43 85.16
Table 6: Impact of rDPO objective across general and hallucination benchmarks for VLMs, and comparisons with baselines.
Results
5 Discussions
Role of Image $v_{l}$
$v_{l}$ is one of the top-10 retrieved images corresponding to the original image $v$, and qualitatively, the images $v$ and $v_{l}$ are semantically similar in terms of scenes, objects, and composition. This retrieval strategy is intended to ensure that $v_{l}$ shares sufficient visual context with $v$, making it a plausible alternative grounding for the instruction $x$. Furthermore, we compute the cosine similarity between the CLIP embeddings of the caption of $v$ (by prompting "Describe this image in detail.") and three types of images: the original image $v$, a retrieved image $v_{l}$, and a randomly selected image $v_{r}$. The average cosine similarities are $0.2780,0.2382,0.0688$, respectively, which indicates that $v_{l}$ retains significant semantic similarity with $v$ and is far more aligned than an unrelated image $v_{r}$. Based on this, we interpret $v_{l}$ as a “rejected input image” to the original instruction $x$: it provides a visually plausible but suboptimal context, under which the response $y_{w}$ should be less preferred compared to when conditioned on $v$.
Discussion with mDPO
$\mathcal{L}_{\text{CoDPO}}=-\mathbb{E}_{(x,v,y_{w},y_{l})\sim\mathcal{D}}$
In contrast, Re-Align extends and enhances this approach by incorporating a more semantically meaningful visual preference pair. Instead of relying solely on random crops, Re-Align retrieves a relevant image from the same dataset that corresponds to the original input. This retrieval-based augmentation provides a stronger contrastive signal, improving the model’s ability to discern fine-grained visual details and reducing spurious correlations. Moreover, beyond mitigating hallucinations in VLMs, Re-Align has been demonstrated that it also significantly enhance performance on general VQA tasks.
6 Related Work
Reinforcement Learning from Human Feedback
Vision Language Models
Large Vision Language Models (VLMs) (Li et al., 2022, 2023a; Liu et al., 2024a; Li et al., 2024b; Meta, 2024; Bai et al., 2023; Wang et al., 2024b; Lu et al., 2024; Wu et al., 2024) extended the understanding and reasoning capabilities of Large Language Models (LLMs) (Devlin et al., 2018; Radford et al., 2019; Brown et al., 2020; Team et al., 2023; Roziere et al., 2023; Touvron et al., 2023a, b; Raffel et al., 2020; Yang et al., 2024; Team, 2024; Pan et al., 2023) into the visual domain. By integrating vision encoders, such as CLIP (Radford et al., 2021b), image patches are first converted into embeddings and then projected to align with text embedding space, unlocking unprecedented cross-modal applications in the real world, such as biomedical imaging (Moor et al., 2023; Li et al., 2024a), autonomous systems (Shao et al., 2024; Tian et al., 2024; Sima et al., 2023; Xing et al., 2024b; Ma et al., 2025; Wang et al., 2025), and robotics (Rana et al., 2023; Kim et al., 2024; Xing et al., 2025).
Alignment of Vision Language Models
Current VLMs often suffer from hallucinations, producing inaccurate or misleading information that fails to accurately represent the content of the provided image (Zhu et al., 2024; Bai et al., 2024; Qian et al., 2025). Such misalignments can have catastrophic consequences when these models are deployed in real-world scenarios (Xing et al., 2024a). To address cross-modality hallucinations, recent research has primarily focused on applying direct preference optimization (Deng et al., 2024; Zhou et al., 2024a; Fang et al., 2024; Zhou et al., 2024b; Guo et al., 2024; Chen et al., 2024; Wang et al., 2024c; Yu et al., 2024b; Li et al., 2023b; Wang et al., 2024a) or contrastive learning (Sarkar et al., 2024) on the curated datasets with preference signals, and utilizing model editing techniques (Liu et al., 2024b; Yu et al., 2024a).
7 Conclusion
In this paper, a novel framework, Re-Align, for aligning VLMs to mitigate hallucinations is proposed. Our approach leverages image retrieval to deliberately induce segment-level hallucinations, thereby generating plausible and natural preference signals. By integrating the retrieved images, a dual-preference dataset that encompasses both textual and visual cues is curated. Furthermore, we propose the rDPO objective, an extension of DPO that includes an additional visual preference optimization objective, to enhance the alignment process with valuable visual preference signals. Comprehensive empirical results from a range of general VQA and hallucination benchmarks demonstrate that Re-Align effectively reduces hallucinations in VLMs while enhancing their overall performance. Moreover, it demonstrates superior scalability across various model architectures and sizes.
Limitations
Although Re-Align has demonstrated superior performance on both hallucination and general VQA benchmarks, it does not always achieve state-of-the-art results on general tasks; in some cases, its performance is even worse than that of vanilla VLMs. Future research could explore strategies to eliminate this alignment tax or identify an optimal balance for this trade-off. The potential risks of this work align with the general challenges of RLHF alignment. As more powerful alignment techniques are developed, they may inadvertently empower adversarial approaches that exploit these models, potentially leading to unfair or discriminatory outputs. Meanwhile, these adversarial strategies can be used to generate negative samples, which can ultimately contribute to the development of more robust and reliable VLMs over time.
Appendix A Overview of Re-Align